# Demystifying Spectral Feature Learning for Instrumental Variable Regression

### v2

@@ Appendix D Estimating spectral alignment
$C_{Z}^{1/2}C_{X}^{1/2}$ is a $d\times d$ matrix, let its SVD be $C_{Z}^{1/2}C_{X}^{1/2}=O\Sigma_{\varphi,\psi}P^{*}$, where $\Sigma_{\varphi,\psi}$ is a positive-definite diagonal matrix with diagonal entries $\sigma_{\varphi,\psi,i}$. Now we can write
$\widehat{\mathcal{T}}_{d}=(\Psi^{\prime}O)\Sigma_{\varphi,\psi}(\Phi^{\prime}P)^{*}.$
Finally letting $u_{\varphi,\psi,i}=\Psi^{\prime}Oe_{i}$, and $v_{\varphi,\psi,i}=\Phi^{\prime}Pe_{i}$, with the corresponding operators $U_{\varphi,\psi}\colon\mathbb{R}^{d}\to L^{2}(Z)$, and $V_{\varphi,\psi}\colon\mathbb{R}^{d}\to L^{2}(X)$, we get the SVD
$\widehat{\mathcal{T}}_{d}=U_{\varphi,\psi}\Sigma_{\varphi,\psi}V_{\varphi,\psi}^{*}.$
@@
D.0.0.0.1 Rationale for the alignment approximation method.
We are interested in evaluating the length of the projection of $h_{0}$ onto the leading learned singular functions of $\widehat{\mathcal{T}}_{d}$, that is of $\langle v_{\varphi,\psi,i},h_{0}\rangle$. Following Eq.˜15, we have
$\langle v_{\varphi,\psi,i},h_{0}\rangle=\sigma_{\varphi,\psi,i}^{-1}\langle\widehat{\mathcal{T}}_{d}^{*}u_{\varphi,\psi,i},h_{0}\rangle=\sigma_{\varphi,\psi,i}^{-1}\langle u_{\varphi,\psi,i},\widehat{\mathcal{T}}_{d}h_{0}\rangle=\sigma_{\varphi,\psi,i}^{-1}\langle u_{\varphi,\psi,i},\mathcal{T}h_{0}\rangle+\sigma_{\varphi,\psi,i}^{-1}\langle u_{\varphi,\psi,i},(\widehat{\mathcal{T}}_{d}-\mathcal{T})h_{0}\rangle.$
The first term on the RHS above can be equivalently written as $\sigma_{\varphi,\psi,i}^{-1}\langle u_{\varphi,\psi,i},\mathcal{T}h_{0}\rangle=\sigma_{\varphi,\psi,i}^{-1}\mathbb{E}[u_{\varphi,\psi,i}(Z)Y]$, for which it is easy to compute a sample-based approximation. This approximation should be close to the true value of $\langle v_{\varphi,\psi,i},h_{0}\rangle$ provided that the second RHS term is small. Letting $\mathcal{T}_{d}$ be the rank-$d$ SVD truncation of $\mathcal{T}$ we further telescope the second term on the RHS to get
$\sigma_{\varphi,\psi,i}^{-1}\left(\langle u_{\varphi,\psi,i},(\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d})h_{0}\rangle+\langle u_{\varphi,\psi,i},(\mathcal{T}_{d}-\mathcal{T})h_{0}\rangle\right).$ (16)
Since $\widehat{\mathcal{T}}_{d}$ should converge to $\mathcal{T}_{d}$ for a sufficiently flexible feature-learning model class, it is sensible to bound the norm of the first term above with $\sigma_{\varphi,\psi,i}^{-1}\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|\|h_{0}\|$. For any $i\leq d$, this upper bound should converge to 0 with the size of the feature-learning sample size. To upper-bound the second term of Eq.˜16, note that that by Wedin Sin-$\Theta$ Theorem, we have
$\|\Pi_{u_{1:d,\varphi,\psi}}-\Pi_{u_{1:d}}\|\leq\frac{\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|}{\sigma_{d}}.$
@@
$\left|\langle u_{\varphi,\psi,i},(\mathcal{T}_{d}-\mathcal{T})h_{0}\rangle\right|\leq\|\Pi_{u_{\varphi,\psi,1:d}}(\mathcal{T}_{d}-\mathcal{T})h_{0}\|\leq\underbrace{\|\Pi_{u_{1:d}}(\mathcal{T}_{d}-\mathcal{T})h_{0}\|}_{=0}+\frac{\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|}{\sigma_{d}}\cdot\|(\mathcal{T}_{d}-\mathcal{T})h_{0}\|.$
@@
$\left|\langle v_{\varphi,\psi,i},h_{0}\rangle-\sigma_{\varphi,\psi,i}^{-1}\mathbb{E}[Yu_{\varphi,\psi,i}]\right|\leq\sigma_{\varphi,\psi,i}^{-1}\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|\|h_{0}\|+\sigma_{d}^{-1}\sigma_{\varphi,\psi,i}^{-1}\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|\|(\mathcal{T}_{d}-\mathcal{T})h_{0}\|.$ (17)
The essential takeaway from the bound above is that as long as the feature-learning sample size can increase to ensure $\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|$ is small, approximating the alignment with $\sigma_{\varphi,\psi,i}^{-1}\mathbb{E}[Yu_{\varphi,\psi,i}]$ is consistent. Moreover, the bound becomes tighter when,
@@
    $\sigma_{\varphi,\psi,i}$ is big (close to 1), which corresponds to a slow spectral decay or looking at the top singular function.
@@
We can additionally compare how close is $\sigma_{\varphi,\psi,i}^{-1}\mathbb{E}[Yu_{\varphi,\psi,i}]$ to the alignment with the true eigenfunction of $\mathcal{T}$: $\langle v_{i},h_{0}\rangle$, so that we can evaluate if we are in the “good” or “bad” scenario. For $i\leq d$, denote $\alpha_{i}=\langle v_{\varphi,\psi,i},h_{0}\rangle$ and $\hat{\alpha}_{i}=\sigma_{\varphi,\psi,i}^{-1}\mathbb{E}[Yu_{\varphi,\psi,i}(Z)]$. The previous bounds shows that
$|\alpha_{i}-\hat{\alpha}_{i}|\leq\text{Err}_{i},\qquad\text{Err}_{i}=\sigma_{\varphi,\psi,i}^{-1}\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|\|h_{0}\|+\sigma_{d}^{-1}\sigma_{\varphi,\psi,i}^{-1}\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|\|(\mathcal{T}_{d}-\mathcal{T})h_{0}\|.$
@@
$\left|\|\Pi_{v_{\varphi,\psi,1:d}}h_{0}\|^{2}-\sum_{i=1}^{d}\sigma_{\varphi,\psi,i}^{-2}|\mathbb{E}[Yu_{\varphi,\psi,i}(Z)]|^{2}\right|=\left|\sum_{i=1}^{d}(\alpha_{i}^{2}-\hat{\alpha}_{i}^{2})\right|=\left|\sum_{i=1}^{d}(\alpha_{i}-\hat{\alpha}_{i})(\alpha_{i}+\hat{\alpha}_{i})\right|\leq\sum_{i=1}^{d}|\alpha_{i}-\hat{\alpha}_{i}|\cdot|\alpha_{i}+\hat{\alpha}_{i}|$
@@
$\left|\|\Pi_{v_{\varphi,\psi,1:d}}h_{0}\|^{2}-\sum_{i=1}^{d}\sigma_{\varphi,\psi,i}^{-2}|\mathbb{E}[Yu_{\varphi,\psi,i}(Z)]|^{2}\right|\leq\sum_{i=1}^{d}\text{Err}_{i}\cdot(2\|h_{0}\|+\text{Err}_{i})=2\|h_{0}\|\sum_{i=1}^{d}\text{Err}_{i}+\sum_{i=1}^{d}(\text{Err}_{i})^{2}$
@@
Finally, to assess that we get the alignment to the eigenfunctions of $\mathcal{T}$ observe that
$\left|\|\Pi_{v_{\varphi,\psi,1:d}}h_{0}\|-\|\Pi_{v_{1:d}}h_{0}\|\right|\leq\left\|\left(\Pi_{u_{\varphi,\psi,1:d}}-\Pi_{u_{1:d}}\right)h_{0}\right\|\leq\frac{\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|}{\sigma_{d}}\|h_{0}\|.$
D.0.0.0.2 Practical considerations.
@@
In practice, we want to perform sample-based approximations of the procedure described above, and while the top singular functions of $\mathcal{T}$ and the corresponding singular values can be reliably estimated, learning the bottom of the spectrum is more unreliable. Given that we divide $\mathbb{E}[Yu_{\varphi,\psi,i}(Z)]$ by $\sigma_{i,\varphi,\psi}$, for singular values close to 0, a small error in estimating them is inflated by the inversion. Hence, we resort to a heuristic that allows us to decide which features and singular functions are learned reliably. Let $\widehat{\mathcal{T}}_{d}=\widehat{U}_{\varphi,\psi}\widehat{\Sigma}_{\varphi,\psi}\widehat{V}_{\varphi,\psi}$ be a finite-sample approximation of the SVD of $\widehat{\mathcal{T}}_{d}$. That is, we perform the SVD computation procedures of the preceding paragraph but relying on sample feature covariance matrices $\widehat{C}_{X}$ and $\widehat{C}_{Z}$. After fitting the SVD once, we recompute the covariance of $\widehat{u}_{\varphi,\psi,i}$ and $\widehat{v}_{\varphi,\psi,i}$ on resampled $(Z^{(b)},X^{(b)})$ data with b=1,,Bb=1,...,B and extract its diagonal terms, which we refer to as $\hat{\sigma}_{\varphi,\psi,1:d}^{(b)}$. Letting the original singular value estimates be $\hat{\sigma}_{\varphi,\psi,i}\doteq\hat{\sigma}_{\varphi,\psi,i}^{(0)}$, we can use the sets $(\hat{\sigma}_{\varphi,\psi,i}^{(b)})_{b=0}^{B}$ to evaluate the reliability of a given singular value estimate. If the variance of the estimates in a given collection of $\hat{\sigma}_{\varphi,\psi,k}^{(b)}$ is large relative to their mean, we treat the value as unreliable and discard all the singular values $\sigma_{\varphi,\psi,i\geq k}$ from further analysis.
@@
For the remaining singular values, $\hat{\sigma}_{\varphi,\psi,1},\cdots,\hat{\sigma}_{\varphi,\psi,k}$, which we treat as “reliably learned” we utilise the quantiles of bootstrap samples to construct pseudo-confidence-sets $\hat{\sigma}_{\varphi,\psi,i}^{\min},\hat{\sigma}_{\varphi,\psi,i}^{\max}]$ for $\sigma_{\varphi,\psi,i}$. We use these as proxies for upper and lower bounds on $\sigma_{\varphi,\psi,i}$. These provide us with matching pseudo-bounds on our estimates of $\langle v_{\varphi,\psi,i},h_{0}\rangle$. That is, the “confidence set” for $\langle v_{i},h_{0}\rangle$ is the interval
$\left[(\hat{\sigma}_{\varphi,\psi,i}^{\max})^{-1}\widehat{\mathbb{E}}[Y\hat{u}_{\varphi,\psi,i}(Z)],(\hat{\sigma}_{\varphi,\psi,i}^{\min})^{-1}\widehat{\mathbb{E}}[Y\hat{u}_{\varphi,\psi,i}(Z)]\right],$
together with the central estimate $(\hat{\sigma}_{\varphi,\psi,i}^{(0)})^{-1}\widehat{\mathbb{E}}[Y\hat{u}_{\varphi,\psi,i}(Z)]$.
D.1 Spectral alignment in synthetic examples
@@
Note that whenever the structural function is available, as is the case in our one-dimensional synthetic example and in the dSprites experiment, one can evaluate the correctness of the approximation proposed above by comparing it to $\langle h_{0},v_{\varphi,\psi,i}\rangle$. In Figure 5, we observe that the method is reliable if the decay of singular values is sufficiently slow for feature-learning to pick up on the singular function pairs reliably. For $c_{\sigma}=0.8$ the approximations are very good while for $c_{\sigma}=0.2$ they become entirely unreliable. A more thorough theoretical analysis of these estimates is a topic warranting further investigation but these observations are generally in line with Eq.˜17. In particular, small values of $\sigma_{\varphi,\psi,i}$ inflate the approximation error.
@@
D.2 dSprites is in the good regime
We fitted the spectral-feature models for dSprites using the same architectures as proposed by [34]. Analogously to the fully synthetic experiment, we are able to evaluate the spectrum of $\widehat{\mathcal{T}}$ and directly measure where in it, $h_{0}$ is supported. As seen in Figure˜6, the top of the learned spectrum is very flat, and hence, in line with what we observed in the previous example, the alignment estimates are reliable. Our experiments confirm that $h_{0}$ is usually spanned by the leading 32 spectral features and 85% of its squared norm spanned by the leading 10 functions. Hence, dSprites is indeed an example of the good case.
(a) Distributions of resampled values of $\hat{\sigma}_{\varphi,\psi,i}^{(b)}$ evaluated for a single $\widehat{\mathcal{T}}_{d}$ model.

### v3

@@ Appendix F Estimating spectral alignment
$C_{Z}^{1/2}C_{X}^{1/2}$ is a $d\times d$ matrix, let its SVD be $C_{Z}^{1/2}C_{X}^{1/2}=O\Sigma_{\phi,\psi}P^{*}$, where $\Sigma_{\phi,\psi}$ is a positive-definite diagonal matrix with diagonal entries $\sigma_{\phi,\psi,i}$. Now we can write
$\widehat{\mathcal{T}}_{d}=(\Psi^{\prime}O)\Sigma_{\phi,\psi}(\Phi^{\prime}P)^{*}.$
Finally letting $u_{\phi,\psi,i}=\Psi^{\prime}Oe_{i}$, and $v_{\phi,\psi,i}=\Phi^{\prime}Pe_{i}$, with the corresponding operators $U_{\phi,\psi}\colon\mathbb{R}^{d}\to L^{2}(Z)$, and $V_{\phi,\psi}\colon\mathbb{R}^{d}\to L^{2}(X)$, we get the SVD
$\widehat{\mathcal{T}}_{d}=U_{\phi,\psi}\Sigma_{\phi,\psi}V_{\phi,\psi}^{*}.$
@@
Rationale for the alignment approximation method.
We are interested in evaluating the length of the projection of $h_{0}$ onto the leading learned singular functions of $\widehat{\mathcal{T}}_{d}$, that is of $\langle v_{\phi,\psi,i},h_{0}\rangle$. Following Eq.˜10, we have
$\langle v_{\phi,\psi,i},h_{0}\rangle$ $=\sigma_{\phi,\psi,i}^{-1}\langle\widehat{\mathcal{T}}_{d}^{*}u_{\phi,\psi,i},h_{0}\rangle=\sigma_{\phi,\psi,i}^{-1}\langle u_{\phi,\psi,i},\widehat{\mathcal{T}}_{d}h_{0}\rangle$
$=\sigma_{\phi,\psi,i}^{-1}\langle u_{\phi,\psi,i},\mathcal{T}h_{0}\rangle+\sigma_{\phi,\psi,i}^{-1}\langle u_{\phi,\psi,i},(\widehat{\mathcal{T}}_{d}-\mathcal{T})h_{0}\rangle.$
The first term on the RHS above can be equivalently written as $\sigma_{\phi,\psi,i}^{-1}\langle u_{\phi,\psi,i},\mathcal{T}h_{0}\rangle=\sigma_{\phi,\psi,i}^{-1}\mathbb{E}[u_{\phi,\psi,i}(Z)Y]$, for which it is easy to compute a sample-based approximation. This approximation should be close to the true value of $\langle v_{\phi,\psi,i},h_{0}\rangle$ provided that the second RHS term is small. Letting $\mathcal{T}_{d}$ be the rank-$d$ SVD truncation of $\mathcal{T}$ we further telescope the second term on the RHS to get
$\sigma_{\phi,\psi,i}^{-1}\left(\langle u_{\phi,\psi,i},(\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d})h_{0}\rangle+\langle u_{\phi,\psi,i},(\mathcal{T}_{d}-\mathcal{T})h_{0}\rangle\right).$ (17)
Since $\widehat{\mathcal{T}}_{d}$ should converge to $\mathcal{T}_{d}$ for a sufficiently flexible feature-learning model class, it is sensible to bound the norm of the first term above with $\sigma_{\phi,\psi,i}^{-1}\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|\|h_{0}\|$. For any $i\leq d$, this upper bound should converge to 0 with the size of the feature-learning sample size. To upper-bound the second term of Eq.˜17, note that by Wedin Sin-$\Theta$ Theorem, we have
$\|\Pi_{u_{1:d,\phi,\psi}}-\Pi_{u_{1:d}}\|\leq\frac{\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|}{\sigma_{d}}.$
@@
$\left|\langle u_{\phi,\psi,i},(\mathcal{T}_{d}-\mathcal{T})h_{0}\rangle\right|\leq\|\Pi_{u_{\phi,\psi,1:d}}(\mathcal{T}_{d}-\mathcal{T})h_{0}\|\leq\underbrace{\|\Pi_{u_{1:d}}(\mathcal{T}_{d}-\mathcal{T})h_{0}\|}_{=0}+\frac{\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|}{\sigma_{d}}\cdot\|(\mathcal{T}_{d}-\mathcal{T})h_{0}\|.$
@@
$\left|\langle v_{\phi,\psi,i},h_{0}\rangle-\sigma_{\phi,\psi,i}^{-1}\mathbb{E}[Yu_{\phi,\psi,i}]\right|\leq\sigma_{\phi,\psi,i}^{-1}\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|\|h_{0}\|+\sigma_{d}^{-1}\sigma_{\phi,\psi,i}^{-1}\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|\|(\mathcal{T}_{d}-\mathcal{T})h_{0}\|.$ (18)
The essential takeaway from the bound above is that as long as the feature-learning sample size can increase to ensure $\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|$ is small, approximating the alignment with $\sigma_{\phi,\psi,i}^{-1}\mathbb{E}[Yu_{\phi,\psi,i}]$ is consistent. Moreover, the bound becomes tighter when,
@@
    $\sigma_{\phi,\psi,i}$ is big (close to 1), which corresponds to a slow spectral decay or looking at the top singular function.
@@
We can additionally compare how close is $\sigma_{\phi,\psi,i}^{-1}\mathbb{E}[Yu_{\phi,\psi,i}]$ to the alignment with the true eigenfunction of $\mathcal{T}$: $\langle v_{i},h_{0}\rangle$, so that we can evaluate if we are in the “good” or “bad” scenario. For $i\leq d$, denote $\alpha_{i}=\langle v_{\phi,\psi,i},h_{0}\rangle$ and $\hat{\alpha}_{i}=\sigma_{\phi,\psi,i}^{-1}\mathbb{E}[Yu_{\phi,\psi,i}(Z)]$. The previous bound shows that
$|\alpha_{i}-\hat{\alpha}_{i}|\leq\text{Err}_{i},\qquad\text{Err}_{i}=\sigma_{\phi,\psi,i}^{-1}\|\widehat{\mathcal{T}}_{d}-\mathcal{T}_{d}\|\|h_{0}\|+\sigma_{d}^{-1}\sigma_{\phi,\psi,i}^{-1}\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|\|(\mathcal{T}_{d}-\mathcal{T})h_{0}\|.$
@@
$\left|\|\Pi_{v_{\phi,\psi,1:d}}h_{0}\|^{2}-\sum_{i=1}^{d}\sigma_{\phi,\psi,i}^{-2}|\mathbb{E}[Yu_{\phi,\psi,i}(Z)]|^{2}\right|$ $=\left|\sum_{i=1}^{d}(\alpha_{i}^{2}-\hat{\alpha}_{i}^{2})\right|=\left|\sum_{i=1}^{d}(\alpha_{i}-\hat{\alpha}_{i})(\alpha_{i}+\hat{\alpha}_{i})\right|$
$\leq\sum_{i=1}^{d}|\alpha_{i}-\hat{\alpha}_{i}|\cdot|\alpha_{i}+\hat{\alpha}_{i}|$
@@
$\left|\|\Pi_{v_{\phi,\psi,1:d}}h_{0}\|^{2}-\sum_{i=1}^{d}\sigma_{\phi,\psi,i}^{-2}|\mathbb{E}[Yu_{\phi,\psi,i}(Z)]|^{2}\right|$ $\leq\sum_{i=1}^{d}\text{Err}_{i}\cdot(2\|h_{0}\|+\text{Err}_{i})$
$=2\|h_{0}\|\sum_{i=1}^{d}\text{Err}_{i}+\sum_{i=1}^{d}(\text{Err}_{i})^{2}$
@@
Finally, to assess how well we get the alignment to the eigenfunctions of $\mathcal{T}$ observe that
$\left|\|\Pi_{v_{\phi,\psi,1:d}}h_{0}\|-\|\Pi_{v_{1:d}}h_{0}\|\right|\leq\left\|\left(\Pi_{v_{\phi,\psi,1:d}}-\Pi_{v_{1:d}}\right)h_{0}\right\|\leq\frac{\|\mathcal{T}_{d}-\widehat{\mathcal{T}}_{d}\|}{\sigma_{d}}\|h_{0}\|.$
Practical considerations.
@@
In practice, we want to perform sample-based approximations of the procedure described above, and while the top singular functions of $\mathcal{T}$ and the corresponding singular values can be reliably estimated, learning the bottom of the spectrum is more unreliable. Given that we divide $\mathbb{E}[Yu_{\phi,\psi,i}(Z)]$ by $\sigma_{i,\phi,\psi}$, for singular values close to 0, a small error in estimating them is inflated by the inversion. Hence, we resort to a heuristic that allows us to decide which features and singular functions are learned reliably. Let $\widehat{\mathcal{T}}_{d}=\widehat{U}_{\phi,\psi}\widehat{\Sigma}_{\phi,\psi}\widehat{V}_{\phi,\psi}$ be a finite-sample approximation of the SVD of $\widehat{\mathcal{T}}_{d}^{*}$. That is, we perform the SVD computation procedures of the preceding paragraph but relying on sample feature covariance matrices $\widehat{C}_{X}$ and $\widehat{C}_{Z}$. After fitting the SVD once, we recompute the covariance of $\widehat{u}_{\phi,\psi,i}$ and $\widehat{v}_{\phi,\psi,i}$ on resampled $(Z^{(b)},X^{(b)})$ data with b=1,,Bb=1,...,B and extract its diagonal terms, which we refer to as $\hat{\sigma}_{\phi,\psi,1:d}^{(b)}$. Letting the original singular value estimates be $\hat{\sigma}_{\phi,\psi,i}\doteq\hat{\sigma}_{\phi,\psi,i}^{(0)}$, we can use the sets $(\hat{\sigma}_{\phi,\psi,i}^{(b)})_{b=0}^{B}$ to evaluate the reliability of a given singular value estimate. If the variance of the estimates in a given collection of $\hat{\sigma}_{\phi,\psi,k}^{(b)}$ is large relative to their mean, we treat the value as unreliable and discard all the singular values $\sigma_{\phi,\psi,i\geq k}$ from further analysis.
@@
For the remaining singular values, $\hat{\sigma}_{\phi,\psi,1},\cdots,\hat{\sigma}_{\phi,\psi,k}$, which we treat as “reliably learned” we utilise the quantiles of bootstrap samples to construct pseudo-confidence-sets $\hat{\sigma}_{\phi,\psi,i}^{\min},\hat{\sigma}_{\phi,\psi,i}^{\max}]$ for $\sigma_{\phi,\psi,i}$. We use these as proxies for upper and lower bounds on $\sigma_{\phi,\psi,i}$. These provide us with matching pseudo-bounds on our estimates of $\langle v_{\phi,\psi,i},h_{0}\rangle$. That is, the “confidence set” for $\langle v_{i},h_{0}\rangle$ is the interval
$\left[(\hat{\sigma}_{\phi,\psi,i}^{\max})^{-1}\widehat{\mathbb{E}}[Y\hat{u}_{\phi,\psi,i}(Z)],(\hat{\sigma}_{\phi,\psi,i}^{\min})^{-1}\widehat{\mathbb{E}}[Y\hat{u}_{\phi,\psi,i}(Z)]\right],$
together with the central estimate $(\hat{\sigma}_{\phi,\psi,i}^{(0)})^{-1}\widehat{\mathbb{E}}[Y\hat{u}_{\phi,\psi,i}(Z)]$.
F.1 Spectral alignment in synthetic examples
@@
Note that whenever the structural function is available, as is the case in our one-dimensional synthetic example and in the dSprites experiment, one can evaluate the correctness of the approximation proposed above by comparing it to $\langle h_{0},v_{\phi,\psi,i}\rangle$. In Figure 6, we observe that the method is reliable if the decay of singular values is sufficiently slow for feature-learning to pick up on the singular function pairs reliably. For $c_{\sigma}=0.8$ the approximations are very good while for $c_{\sigma}=0.2$ they become entirely unreliable. A more thorough theoretical analysis of these estimates is a topic warranting further investigation but these observations are generally in line with Eq.˜18. In particular, small values of $\sigma_{\phi,\psi,i}$ inflate the approximation error.
@@
F.2 dSprites is in the good regime
We fitted the spectral-feature models for dSprites using the same architectures as proposed by [36]. Analogously to the fully synthetic experiment, we are able to evaluate the spectrum of $\widehat{\mathcal{T}}$ and directly measure where in it, $h_{0}$ is supported. As seen in Figure˜7, the top of the learned spectrum is very flat, and hence, in line with what we observed in the previous example, the alignment estimates are reliable. Our experiments confirm that $h_{0}$ is usually spanned by the leading 32 spectral features and 85% of its squared norm spanned by the leading 10 functions. Hence, dSprites is indeed an example of the good case.
(a) Distributions of resampled values of $\hat{\sigma}_{\phi,\psi,i}^{(b)}$ evaluated for a single $\widehat{\mathcal{T}}_{d}$ model.